# Sensory Data Sharing in ISAC System: Fundamental Limits and Waveform Design

### v2

@@ Abstract
The simultaneous acquisition and sharing of sensory data through a dual-functional signaling strategy, termed the communication-assisted sensing (CAS) system in this paper, has the potential to provide users with beyond-line-of-sight sensing capabilities. We mainly focus on three primary aspects, namely, the information-theoretic framework, the optimal distribution of channel input, and the optimal waveform design for Gaussian signals. First, we establish the information-theoretic framework and develop a modified source-channel separation theorem (MSST) tailored for CAS systems. The proposed MSST elucidates the relationship between achievable distortion, coding rate, and communication channel capacity in cases where the distortion metric is separable for sensing and communication (S&C) processes. Second, we present an optimal channel input design for dual-functional signaling, which aims to minimize CAS distortion under the constraints of the MSST and resource budget. We then conceive a two-step Blahut-Arimoto (BA)-based optimal search algorithm to numerically solve the functional optimization problem. Third, in light of the current signaling strategy, we further propose an optimal waveform design for Gaussian signaling in multi-input multi-output (MIMO) CAS systems. The associated covariance matrix optimization problem is addressed using a successive convex approximation (SCA)-based waveform design algorithm. Finally, we provide numerical simulation results to demonstrate the effectiveness of the proposed algorithms and to show the unique performance tradeoff between S&C processes.
Index Terms:
Integrated sensing and communication, communication assisted sensing, rate-distortion theory, source-channel separation theorem, Blahut-Arimoto algorithm.
@@ I Introduction
Integrated sensing and communication (ISAC) systems are widely acknowledged for their potential to enhance sensing and communications (S&C) performance by sharing hardware and spectrum resources [2, 3]. Over the past few decades, substantial research efforts have focused on improving resource efficiency and mitigating mutual interference through waveform design [4], beamforming [5], and dedicated signal processing techniques [6]. To further explore the coordination gains between S&C, the authors of [7] proposed a sensing-assisted communication scheme, which leverages sensing capability of the transmitting ISAC signal, such as beam tracking and prediction, to establish user links thereby reducing the communication overhead. On the other hand, sensing-as-a-service is expected to become a key feature of upcoming 6G perceptive networks [8], which facilitates the reconsideration of how to share sensory measurement or target-related information (such as state variables) among devices or a fusion center within ISAC systems.
Figure 1: The application scenarios of the CAS system.
I-A Communication-Assisted Sensing Systems
The sharing of sensory data to extend coverage and enhance sensing accuracy has garnered extensive attention in traditional networks. For instance, the authors of [9] proposed a beyond-line-of-sight (BLoS) sensing scheme that shares measurements collected by vehicle-mounted sensors among users, enabling the detection of obstructed or distant targets. However, in ISAC systems, sensory data acquisition and sharing can be achieved simultaneously using a dual-functional signaling strategy, eliminating the need for additional sensors. This approach, referred to as the communication-assisted sensing (CAS) system in [10, 11], differs from traditional methods in that BLoS sensing is an intrinsic capability of the network.
Let us take the scenario in Fig. 1 as an example to illustrate the unique challenges in CAS sytems. The base station (BS) or roadside unit (RSU) with favorable visibility illuminates targets and captures observations during the sensing process while simultaneously transmitting the acquired sensory information to the end-users during the communication process. In a dual-functional signaling strategy, the ISAC waveform is used both to sense the current targets’ states and to convey the previously acquired state information to the end-users. This strategy inherently couples the S&C processes, giving rise to two unresolved issues in previous works.
1) The information theoretical framework of the CAS system has not been validated from a coding theory perspective, remaining an unclear operational meaning.
2) The ISAC waveform is assumed to be a Gaussian distribution in [11], which may not be necessarily optimal for the CAS scenario. Note the fact that Gaussian distribution is communication-optimal under Gaussian channels, whereas 2-ary pulse amplitude modulation is sensing-optimal [12].
@@ I-A Communication-Assisted Sensing Systems
This motivates us to delve into the fundamental limits and the optimal waveform design of the dual-functional signaling strategy for CAS systems.
@@
Figure 2: The comparisons of the information-theoretic frameworks between the remote estimation, remote source coding, and the proposed CAS system.
I-B Related Works
@@
Exploring the fundamental limits plays an important role in the ISAC systems [13]. The groundbreaking studies [12, 14] have considered a scenario where the transmitter (Tx) communicates with a user through a memoryless state-dependent channel while simultaneously estimating the state from generalized feedback. The capacity-distortion-cost tradeoff of this channel is characterized to illustrate the optimal achievable rate for reliable communication while maintaining a preset state estimation distortion. The work in [15] reveals the deterministic-random tradeoff between S&C within the dual-functional signaling strategy by characterizing the Cramér-Rao bound (CRB)-communication rate region. Unfortunately, the fundamental limits of CAS systems, whose setups significantly differ from the existing literature, remain widely unexplored.
@@
The information-theoretic framework of the dual-functional signaling strategy for CAS systems is depicted in Fig 2, where $\mathsf{S}$, $\mathsf{Z}$, $\mathsf{X}$, $\mathsf{Y}$, and $\mathsf{\hat{S}}$ represent the source, the sensing channel observation, S&C channels input, the communication channel output, and the recovery at the receiver, respectively. Essentially, the CAS framework shares a similar methodology with conventional remote source coding (RSC), remote estimation (RE), and information bottleneck (IB) problems, by striving to balance the tradeoff between compression complexity (from $\mathsf{Z}$ to $\mathsf{\hat{S}}$) and the informative preservation (from $\mathsf{S}$ to $\mathsf{\hat{S}}$). However, we highlight the unique challenge introduced by adopting the ISAC signal $\mathsf{X}$ as both the S&C channels input.
@@
$\bullet$ Remote source coding: In the RSC scheme, the noisy observation $\mathsf{Z}$ is encoded instead of the original source information $\mathsf{S}$ [16, 17, 18, 19, 20]. Rate-distortion theory is used to quantify the tradeoff between communication rate and achieved distortion, benefiting from the Markov chain $\mathsf{S}\leftrightarrow\mathsf{Z}\leftrightarrow\mathsf{X}\leftrightarrow\mathsf{Y}\leftrightarrow\mathsf{\hat{S}}$ [20]. However, in the CAS system, the observation $\mathsf{Z}$ depends on the ISAC channel input $\mathsf{X}$ due to the dual-functional signaling strategy. This dependency violates the Markov chain assumption, complicating the characterization of the rate-distortion relationship.
$\bullet$ Remote estimation: The RE involves a sensor measuring the state of a linear system and transmitting its observations to a remote estimator over a wireless fading channel [21, 22, 23, 24]. The sensory data-sharing scheme considered in [9] is a typical application in vehicular networks. This widely investigated problem differs from the CAS system in two ways. First, the RE requires additional sensors, whereas sensing capability is an intrinsic function in CAS system. Second, similar to the RSC, the observation $\mathsf{Z}$ and channel input $\mathsf{X}$ are decoupling.
@@
$\bullet$ Information bottleneck: The IB problem is essentially equivalent to the RSC problem in which the distortion metric is specified by the logarithm loss fidelity criterion [25, 26]. For discussion convenience, we focus on the simplified process $\mathsf{S}\leftrightarrow\mathsf{Z}\leftrightarrow\mathsf{\hat{S}}$. The IB problem may be expressed by
@@
where $I(\cdot,\cdot)$ denotes the mutual information (MI) and $\beta$ is the preset positive Lagrange multiplier that balances the tradeoff between compression and relevance. In contrast, when incorporating the ISAC signal $\mathsf{X}$ in CAS system, all variables in the IB problem become coupled. This leads to the following reformulation
[…]
where the tradeoff parameter $\beta$ is replaced by the communication channel capacity $I(\mathsf{X};\mathsf{Y})$ which also depends on $\mathsf{X}$. Evidently, the introduced ISAC signal significantly complicates the IB problem within the context of CAS system.
@@ I-C Our contributions
Considering the aforementioned unique challenges, our aim is to establish the information-theoretic framework for the CAS systems, delving into the fundamental limits and specific waveform design. Compared to the conference version [1], which only discussed the corresponding information-theoretic framework, we further elaborate on the optimal distribution of channel input and propose a novel method for Gaussian waveform design in MIMO CAS systems. For clarity, the main contributions of this work are summarized as follows.
@@
  • First, we establish the information-theoretic framework for CAS systems to illustrate the relationship between achievable distortion, coding rate, and communication channel capacity. We develop a modified source-channel separation theorem (MSST) specific to the cases of separable distortion metric for S&C processes. Compared to the existing works [11, 10], we provide a rigorous proof of the operational meaning for the MSST.
@@
  • Second, we develop a optimal distribution design for the ISAC signal (S&C channel input), where the CAS distortion is minimized while adhering to the MSST and resource constraints. To cope with the functional optimization problem, we conceive a two-step Blahut-Arimoto (BA)-based optimal search algorithm in an effort to tackle the challenges of lacking explicit expressions for the rate-distortion bound and channel capacity.
@@
  • Third, we propose an optimal waveform design scheme for Gaussian signaling in MIMO CAS systems. The associated covariance matrix optimization problem is solved using an successive convex approximation (SCA)-based waveform design algorithm.
@@
  • Finally, we provide numerical simulation results to show the effectiveness of the proposed algorithms, while demonstrating the unique performance tradeoff between S&C in the considered CAS systems.
@@
This paper is structured as follows. We commence with establishing the information-theoretic framework in Section II, including the definitions of distortions, rate-distortion function, and constrained channel capacity. In Section III, we prove the achievability and converse of the proposed MSST. Then, we formulate the optimization problem for ISAC channel input design and develop a two-step BA-based optimal search algorithm in Section IV. In Section V, we present the Gaussian ISAC waveform design method for the MIMO CAS systems. Finally, we provide the simulation results and conclude this paper in Section VI and VII, respectively.
@@
The notations used in this paper are as follows. The uppercase normal letter $\mathsf{A}$, lowercase italic letter $a$, and fraktur letter $\mathcal{A}$ denote a random variable, its realization and a set, respectively. $P_{\mathsf{A}}(a)$ represent a probability distribution function and $Q_{\mathsf{A|B}}(a|b)$ specifies to channel transition probability. Uppercase and lowercase bold letters $\mathbf{A}$ and $\mathbf{a}$ denote the matrix and column vector. $(\cdot)^{T}$, $(\cdot)^{*}$, and $(\cdot)^{H}$ represent the transpose, conjugate, and complex conjugate transpose operations, respectively. $\mathbb{E}\left[\cdot\right]$ is the statistical expectation, and $\text{Tr}\left(\cdot\right)$ is the trace of a matrix.
II Information-theoretic framework of CAS
@@ II-A Sensing and Communication Processes
As shown in Fig. 2, the random variables of the target’s state $\mathsf{S}$, the sensing observation $\mathsf{Z}$, the ISAC signal $\mathsf{X}$, and the communication channel output $\mathsf{Y}$ take values in the sets $\mathcal{S}$, $\mathcal{Z}$, $\mathcal{X}$, and $\mathcal{Y}$, respectively. Here, the state sequence $\{\mathsf{S}_{i}\}_{i\geq 1}$ is independently and identically distributed (i.i.d.) subject to a prior distribution $P_{\mathsf{S}}(s)$. In general, the CAS system may be described as the following S&C processes.
@@
By denoting $R$ as the bit rate, a $(2^{nR},n)$ coding scheme for the CAS systems consists of
@@
The performance of the CAS systems may be evaluated by the distortion between the ground truth and its reconstruction at the receiver. We term this distortion as the CAS distortion, which can be defined by
@@
where $D$ and $B$ represent the CAS distortion and resource budget, respectively. We define the capacity-distortion-cost function $C_{\text{CAS}}(D,B)$ as the infimum of rate $R$ such that the tuple $(R,D,B)$ is achievable. As analyzed in Section I-B, compared to the existing works [12, 14], it is difficult to characterize such a capacity-distortion-cost curve due to the strong coupling between S&C process through the ISAC signal. In this paper, we characterize $C_{\text{CAS}}(D,B)$ within a estimate and compress (EC) strategy which achieves an optimal tradeoff between rate and distortion [27], and leave the general strategy for future research.

### v3

# Simultaneous Sensing Data Acquisition and Sharing in Low-Altitude Wireless Networks: Fundamental Limits and Optimal Signaling

@@ Abstract
In the low-altitude wireless networks, the simultaneous sensing data acquisition and sharing (SDAS) through an ISAC signaling strategy becomes a typical application scenario. In this paper, we mainly investigate three primary aspects of the SDAS system, namely, the information-theoretic framework, the optimal distribution of channel input, and the optimal waveform design for Gaussian signaling. First, we establish the information-theoretic framework and develop a modified source-channel separation theorem (MSST) tailored for the SDAS systems. The proposed MSST elucidates the relationship between achievable distortion, coding rate, and communication channel capacity in cases where the distortion metric is separable for sensing and communication (S&C) processes. Second, we present an optimal channel input design for dual-functional signaling, which aims to minimize SDAS distortion under the constraints of the MSST and resource budget. We then conceive a two-step Blahut-Arimoto (BA)-based optimal search algorithm to numerically solve the functional optimization problem. Third, to provide practical design insights, we further propose an optimal waveform design for Gaussian signaling in multi-input multi-output (MIMO) SDAS systems. The associated covariance matrix optimization problem is addressed using a successive convex approximation (SCA)-based waveform design algorithm. Finally, we provide numerical simulation results to demonstrate the effectiveness of the proposed algorithms, which characterizes the unique performance tradeoff between S&C processes.
@@ I Introduction
Unmanned aerial vehicles (UAVs) play a pivotal role in low-altitude wireless networks (LAWN), serving as temporary communication base stations (BS) or aerial monitoring platforms by utilizing their high mobility [2]. Recently, the integrated sensing and communication (ISAC) technique has revolutionized UAV systems by offering notable advantages such as mutually beneficial interactions between sensing and communication (S&C). The favorable attributes significantly broaden the potential applications of ISAC-empowered UAV, including smart transportation and agriculture [3, 4]. In many scenarios, UAVs are required for target sensing and subsequently transmitting the observed parameters to remote users, especially when direct line-of-sight (LoS) sensing channels between the users and targets are blocked. Thus, the performance is typically quantified by the error between the ground truth and the user’s recovery. This closely resembles the established Remote Estimation problem [5, 6, 7, 8] when the S&C processes are independent.
However, a novel paradigm, which we term simultaneous sensing data acquisition and sharing (SDAS) in this paper, emerges when UAVs leverage dual-functional ISAC signals to simultaneously sense current target states and convey previously acquired state information to end-users. The S&C processes are deeply coupled through the transmitted signal, leading to fundamental differences in both theoretical model and signal design. In what follows, we discuss the key differences between our work and related works from both an application layer and an information-theoretic perspective.
Figure 1: The use case of SDAS systems in LAMN.
I-A Related Works on Application Layer
I-A1 Traditional ISAC Systems
Most research on ISAC-UAV primarily focus on the joint optimization of trajectory planning, resource allocation and waveform design. These studies often consider three key performance indicators: sensing metrics (e.g., signal to interference and noise ratio (SINR), Cramér-Rao Bound) [9], communication metrics (e.g., achievable rate) [10, 11] and energy consumption [12]. Typically, one of these (or a weighted sum) is chosen as the optimization objective, with the others imposed as constraints. Evidently, S&C performance are characterized through individual and even incompatible evaluation system, e.g., sensing for estimation theory and communication for information theory. However, the SDAS framework is quantified by the error between ground truth and user’s recovery, implicitly reflecting the integrated impact of the S&C processes. To our best knowledge, the SDAS distortion has not been well-defined, and its minimization design remains largely unexplored.
I-A2 Relay Systems
Relay systems primarily function by forwarding the signals from a source transmitter to a destination receiver, typically operating in the following two modes: decode-and-forward the signal or simply amplify-and-forward the signal [13, 14]. Despite its structure is similar to that of SDAS, its main purpose is to transmit communication signals in a cooperative manner without actively performing sensing functions. Furthermore, channel estimation in the relay systems employs either a cooperative training method [15] or dividing the estimation into individual source-relay and relay-destination channels [16]. However, a quantitative analysis of how the S&C coupling process affects the SDAS distortion is still lacking.
I-A3 Communication-assisted Sensing Systems
In previous work [17, 18], the authors established a communication-assisted sensing framework and proposed SDAS distortion minimization-based waveform designs for separated S&C and dual-functional waveform schemes, respectively. However, the information-theoretic framework of SDAS has not been rigorously validated from a coding theory perspective, remaining an unclear operational meaning. Moreover, the reliance on Gaussian distribution for dual-functional ISAC waveform design in [18] poses a potential sub-optimal property for the SDAS scenario, especially considering the fact that Gaussian distribution is communication-optimal under Gaussian channel, whereas constant-modulus modulation is sensing-optimal [19].
@@ I-A3 Communication-assisted Sensing Systems
The above research gap motivates us to delve into the fundamental limits and the optimal waveform design of the dual-functional ISAC signaling strategy for the SDAS systems.
@@
Figure 2: The comparisons of the information-theoretic frameworks between the remote estimation, remote source coding, and the SDAS system.
I-B Related Works on Information-theoretic Framework
@@
The pioneering studies [19, 20] have considered a scenario where the transmitter (Tx) communicates with a user through a memoryless state-dependent channel while simultaneously estimating the state from generalized feedback. The capacity-distortion-cost tradeoff of this channel is characterized to illustrate the optimal achievable rate for reliable communication while maintaining a preset state estimation distortion. The authors of [21] further extend the capacity-distortion function to a framework based on backward simultaneous decoding, which does not need the Wyner-Ziv random binning method. Moreover, the work in [22] reveals the deterministic-random tradeoff between S&C within the dual-functional signaling strategy by characterizing the Cramér-Rao bound (CRB)-communication rate region. Unfortunately, the fundamental limits of SDAS systems, whose setups significantly differ from the existing literature, remain widely unexplored [23].
@@
The information-theoretic framework for SDAS systems is depicted in Fig 2, where $\mathsf{S}$, $\mathsf{Z}$, $\mathsf{X}$, $\mathsf{Y}$, and $\mathsf{\hat{S}}$ represent the source, sensing channel observation, S&C channels input, communication channel output, and recovery at the receiver, respectively. Essentially, the SDAS framework shares a similar methodology with conventional remote estimation (RE), remote source coding (RSC), and information bottleneck (IB) problems, by striving to balance the tradeoff between compression complexity (from $\mathsf{Z}$ to $\mathsf{\hat{S}}$) and the informative preservation (from $\mathsf{S}$ to $\mathsf{\hat{S}}$). However, we highlight the unique challenge introduced by adopting the ISAC signal $\mathsf{X}$ as both the S&C channels input.
I-B1 Remote estimation
The RE framework involves a sensor measuring the state of a linear system and transmitting its observations to a remote estimator over a wireless fading channel [5, 6, 7, 8]. Sensing data acquisition in RE scheme is achieved through additional sensors, resulting that the observation $\mathsf{Z}$ and the communication channel input $\mathsf{X}$ are decoupled. In contrast, the observation $\mathsf{Z}$ depends on the ISAC channel input $\mathsf{X}$ due to the dual-functional signaling strategy in SDAS system.
I-B2 Remote source coding
@@
In RSC framework, the noisy observation $\mathsf{Z}$ is encoded instead of the original source information $\mathsf{S}$ [24, 25, 26, 27, 28]. Rate-distortion theory is used to quantify the tradeoff between communication rate and achievable distortion, benefiting from the Markov chain $\mathsf{S}\leftrightarrow\mathsf{Z}\leftrightarrow\mathsf{X}\leftrightarrow\mathsf{Y}\leftrightarrow\mathsf{\hat{S}}$ [28]. However, the dependency between $\mathsf{Z}$ and $\mathsf{X}$ in SDAS violates the Markov chain assumption, complicating the characterization of the rate-distortion relationship.
I-B3 Information bottleneck
@@
The IB problem is essentially equivalent to the RSC problem in which the distortion metric is specified by the logarithm loss fidelity criterion [29, 30]. For discussion convenience, we focus on the simplified process $\mathsf{S}\leftrightarrow\mathsf{Z}\leftrightarrow\mathsf{\hat{S}}$. The IB problem may be expressed by
@@
where $I(\cdot,\cdot)$ denotes the mutual information (MI) and $\beta$ is the preset positive Lagrange multiplier that balances the tradeoff between compression and relevance. In contrast, when incorporating the ISAC signal $\mathsf{X}$ in SDAS system, all variables in the IB problem become coupled. This leads to the following reformulation
$\mathop{\text{max}}\kern 2.0ptI(\mathsf{S};\mathsf{\hat{S}}(\mathsf{X}))\kern 5.0pt\text{subject to}\kern 5.0ptI(\mathsf{Z}(\mathsf{X});\mathsf{\hat{S}}(\mathsf{X}))\leq I(\mathsf{X};\mathsf{Y}),$
where the tradeoff parameter $\beta$ is replaced by the communication channel capacity $I(\mathsf{X};\mathsf{Y})$ which also depends on $\mathsf{X}$. Evidently, the introduced ISAC signal significantly complicates the IB problem within the context of SDAS system.
@@ I-C Our contributions
Considering the aforementioned unique challenges, our aim is to establish the information-theoretic framework for the SDAS systems, characterize its fundamental limits, and design dedicated ISAC waveform for SDAS transmission. Compared to the conference version [1], which only discussed the information-theoretic framework, we further elaborate on the optimal distribution of channel input and propose a novel method for Gaussian waveform design in MIMO SDAS systems. The main contributions are summarized as follows.
@@
  • First, we establish the information-theoretic framework for SDAS systems to illustrate the relationship between achievable distortion, coding rate, and communication channel capacity. We develop a modified source-channel separation theorem (MSST) specific to the cases of separable distortion metric for S&C processes. Compared to the existing works [18, 17], we provide a rigorous proof of the operational meaning for the MSST.
@@
  • Second, we develop a optimal distribution design for the ISAC signal (i.e., S&C channel input), where the SDAS distortion is minimized while adhering to the MSST and resource budget constraints. To cope with the functional optimization problem, we conceive a two-step Blahut-Arimoto (BA)-based optimal search algorithm in an effort to tackle the challenges of lacking explicit expressions for the rate-distortion bound and channel capacity.
@@
  • Third, we propose an optimal waveform design scheme for Gaussian signaling in MIMO SDAS systems. Under the Gaussian assumption, the explicit expressions of the rate-distortion function and channel capacity are derived. Thus, the optimal waveform design is modeled by a covariance matrix optimization problem solved by using an successive convex approximation (SCA)-based algorithm.
@@
  • Finally, we provide numerical simulation results to show the effectiveness of the proposed algorithms, and demonstrate the performance tradeoff between S&C processes in the SDAS systems.
@@
This paper is structured as follows. We commence with establishing the information-theoretic framework in Section II, including the definitions of distortions, rate-distortion function, and constrained channel capacity. In Section III, we prove the achievability and converse of the MSST. Then, we formulate the optimization problem for ISAC channel input design and develop a two-step BA-based search algorithm in Section IV. In Section V, we present the Gaussian ISAC waveform design method for the MIMO SDAS systems. Finally, we provide the simulation results and conclude this paper in Section VI and VII, respectively.
@@
Notations: The uppercase normal letter $\mathsf{A}$, lowercase italic letter $a$, and fraktur letter $\mathcal{A}$ denote a random variable, its realization and a set, respectively. $P_{\mathsf{A}}(a)$ represent a probability distribution function and $Q_{\mathsf{A|B}}(a|b)$ specifies to channel transition probability. Uppercase and lowercase bold letters $\mathbf{A}$ and $\mathbf{a}$ denote the matrix and column vector. $(\cdot)^{T}$, $(\cdot)^{*}$, and $(\cdot)^{H}$ represent the transpose, conjugate, and complex conjugate transpose operations, respectively. $\mathbb{E}\left[\cdot\right]$ is the statistical expectation, and $\text{Tr}\left(\cdot\right)$ is the trace of a matrix.
II Information-theoretic framework of SDAS
@@ II-A Sensing and Communication Processes
As shown in Fig. 2, the random variables of the target’s state $\mathsf{S}$, the sensing observation $\mathsf{Z}$, the ISAC signal $\mathsf{X}$, and the communication channel output $\mathsf{Y}$ take values in the sets $\mathcal{S}$, $\mathcal{Z}$, $\mathcal{X}$, and $\mathcal{Y}$, respectively. Here, the state sequence $\{\mathsf{S}_{i}\}_{i\geq 1}$ is independently and identically distributed (i.i.d.) subject to a prior distribution $P_{\mathsf{S}}(s)$. In general, the SDAS system may be described as the following S&C processes.
@@
By denoting $R$ as the bit rate, a $(2^{nR},n)$ coding scheme for the SDAS systems consists of
@@
The performance of SDAS systems can be evaluated by the distortion between the ground truth and its reconstruction at the receiver. We term this distortion as the SDAS distortion, which can be defined by
@@
[…]